# Edge Weight Prediction For Category-Agnostic Pose Estimation

### v1

@@ C.1 Framework Overview
We base our method on GraphCape [14], introducing three main architectural modifications:
  • Updating the feature extraction module.
@@
  • Adding a graph-prediction module.
@@
  • Feature Extractor: A pre-trained model extracts features from support and query images, resulting in the query feature map $\hat{F_{q}}\in\mathbb{R}^{hw\times C}$ and support keypoint features $\hat{F_{s}}\in\mathbb{R}^{K\times C}$. Support keypoint features are derived by element-wise multiplication between the support image’s feature map and keypoint masks, created using Gaussian kernels centered at the support keypoints. In multi-shot scenarios (e.g., 5-shot), the average of support keypoint features across images in feature space is taken.
  • Transformer Encoder: The transformer encoder fuses information between support keypoint and query patch features. Support keypoints and query features are concatenated before entering the self-attention layer and separated afterward. The output is the refined query feature map $F_{q}$ and refined support keypoint features $F_{s}$.
  • Similarity-Aware Proposal Generator: GraphCape builds on CapeFormer’s two-stage approach, first generating initial keypoint predictions, which are then refined via a DETR-based transformer decoder. The proposal generator aligns support keypoint features with query features, producing similarity maps from which peaks are selected as similarity-aware proposals. To enhance efficiency and adaptability, a trainable inner-product mechanism [34] is used to explicitly model similarity.
  • Graph Transformer Decoder: A transformer decoder network decodes keypoint locations from the query feature map. Each layer contains self-attention, cross-attention, and feed-forward blocks. In this design, GraphCape replaces the original transformer decoder’s feed-forward network with a GCN-based module to incorporate structural priors. To prevent excessive smoothing—a common issue in deep GCNs that can blur node distinctions and degrade performance—GraphCape adds a linear layer for each node following the GCN layer. Each decoder layer predicts coordinate deltas for prior predictions using an iterative refinement strategy [3, 38, 55]. The updated coordinates are created as follows:
@@
C.2 Markovian Attention Bias
Figure 13: Graph Decoder Prediction Layer. Overview of the Transformer decoder architecture, adapted from the GraphCape design. The decoder consists of self-attention, cross-attention, and a graph-based feed-forward network. We incorporate a Markovian Attention Bias into the self-attention mechanism to encourage structural keypoint interactions. Self-attention facilitates adaptive interactions among support keypoints, while cross-attention extracts localization information from the input features. Finally, the decoder refines keypoint features and outputs location predictions.
In category-agnostic pose estimation, encoding spatial relations between keypoints is beneficial for robust localization, especially when dealing with novel objects [14, 33]. Transformer models naturally have a global receptive field, allowing each node (or token) to attend to all others within a layer. However, this flexibility introduces a challenge: Transformers lack inherent structural constraints, so positional dependencies must be explicitly encoded to reflect local relations. While this is often done in sequence data through absolute or relative positional encodings, graphs pose a different challenge, as nodes are not arranged linearly and connectivity is defined by edges.
Our approach aims to better utilize the structural dependencies between keypoints for CAPE. We build on the foundations set by GraphCape [14], which incorporates GCN layers into the feed-forward layers to propagate structural information. However, the GCN layers used in that model were limited by their fixed, nearest-neighbor receptive field, which restricts the model’s ability to capture more complex or distant connections between keypoints.
Thus, we further integrate the graph-prior into the architecture. We follow Ying et al. [48], adding a bias term based on graph connectivity to the self-attention mechanism in the decoder. Denote $a_{ij}$ as the $(i,j)$-element of the Query-Key product attention matrix $a$, resulting in:
@@ C.2 Markovian Attention Bias
This bias term is highly effective in capturing general structure, boosting the performance of GraphCape by around 0.5%. However, it assumes a discrete distance between nodes and is applicable for unweighted adjacency matrices. Our primary objective is to predict real-valued edges, representing the strength of the structural keypoints connections. Thus, we follow Ma et al. [21] and treat the normalized adjacency matrix $\widetilde{A}$ as a right stochastic matrix. We differentiate their implementation by using the bias attention mechanism instead of relative positional encoding.
A stochastic matrix is commonly used to describe the transitions in a Markov chain, where each element $(\widetilde{A})_{ij}$ represents the probability of moving from one state (or node) $v_{i}$ to another state $v_{j}$ in a single step. Specifically, the entry $(\widetilde{A}^{k})_{ij}$ gives the probability of transitioning from node $v_{i}$ to node $v_{j}$ in exactly $k$-hops. This process allows us to capture both direct and indirect relations between keypoints, enabling the model to consider more distant keypoints that may influence localization. Thus, like Ma et al. [21], we build the following matrix:
@@
where $I$ is the identity matrix and the parameter $K$ controls the maximum number of hops considered. We incorporate this graph characteristic into our model’s attention mechanism to enable more nuanced and flexible structural priors. The complete decoder layer is illustrated in Figure 13. Based on Equations 11 and 12, we use the following bias term:
@@
C.3 Training Scheme
@@
We begin by training the base model and establishing strong foundational features necessary for robust localization. We follow [19] and drop the heatmap loss used in GraphCape [14] as we empirically notice that this loss term doesn’t improve performance. Thus, optimization is done using $\mathcal{L}_{offset}$ which is the $L_{1}$ localization loss as in [14]. Once the model has converged, we freeze the feature extractor module and integrate the skeleton predictor. This stage allows the model to further adapt by incorporating specific structural insights provided by the skeleton predictor. For this phase, we add the adjacency loss, resulting in:
@@
In the final phase, we maintain the frozen feature extractor and freeze the skeleton predictor. Then, we integrate the Markovian bias attention. This final stage allows the model to strengthen the model’s capacity to interpret spatial dependencies between keypoints.

### v2

@@ C.1 Framework Overview
We base our method on GraphCape (Hirschorn and Avidan, 2024), introducing two main architectural modifications:
@@
  • Adding a category-agnostic pose-graph prediction module.
@@
  • Feature Extractor: A pre-trained model extracts features from both support and query images. Support keypoint features are derived by element-wise multiplication between the support image’s feature map and keypoint masks, created using Gaussian kernels centered at the support keypoints. In multi-shot scenarios (e.g., 5-shot), the average of support keypoint features across images in feature space is taken. This results in the query feature map $\hat{F_{q}}\in\mathbb{R}^{hw\times C}$ and support keypoint features $\hat{F_{s}}\in\mathbb{R}^{K\times C}$.
  • Transformer Encoder: The transformer encoder fuses information between support keypoint and query patch features. It comprises three transformer blocks, each with a self-attention layer. Support keypoints and query features are concatenated before entering the self-attention layer and separated afterward. The output is the refined query feature map $F_{q}$ and refined support keypoint features $F_{s}$.
  • Similarity-Aware Proposal Generator: GraphCape builds on CapeFormer’s two-stage approach, first generating initial keypoint predictions, which are then refined via a DETR-based transformer decoder. The proposal generator aligns support keypoint features with query features, producing similarity maps from which peaks are selected as similarity-aware proposals. To enhance efficiency and adaptability, a trainable inner-product mechanism (Shi et al., 2022) is used to explicitly model similarity.
  • Graph Transformer Decoder: A transformer decoder network decodes keypoint locations from the query feature map. It contains three layers, each with self-attention, cross-attention, and feed-forward blocks. In this design, GraphCape replaces the simple MLP in the transformer decoder’s feed-forward network with a GCN-based module to incorporate structural priors directly into the keypoint prediction process. To prevent excessive smoothing—a common issue in deep GCNs that can blur node distinctions and degrade performance—GraphCape adds a linear layer for each node following the GCN layer:
    $F_{s}^{k}=W_{linear}\sigma_{act}(W_{adj}F_{s}^{k}\widetilde{A}_{prior}+W_{self}F_{s}^{k})$ (14)
    Where $W_{i}$ are learnable parameters, $\sigma_{act}$ is an activation function (ReLU), and $\widetilde{A}_{prior}\in\mathbb{R}^{K\times K}$ is the symmetrically normalized adjacency matrix.
    Using an iterative refinement strategy (Cai and Vasconcelos, 2018; Teed and Deng, 2020; Zhu et al., 2020), each decoder layer predicts coordinate deltas for prior predictions. This means that the updated coordinates are created as follows:
@@
C.2 Category-Agnostic Pose-Graph Prediction
Figure 15: Category-Agnostic Pose-Graph Predictor. Our pose-graph predictor (as referenced in Figure 3) uses the prior input graph $A_{prior}$, the support image features $F_{s}$ and keypoint features $F_{s}^{k}$ to produce structure-aware keypoint features $F_{refined}^{k}$. Cosine-similarity is then applied to predict the residual adjacency output $\Delta A$.
The architecture of our pose-graph prediction network $f_{\theta}$ is illustrated in Figure 15. Formally, we define a learnable function $f_{\theta}$ to learn the residual pose-graph in the form of an adjacency matrix $\Delta A\in\mathbb{R}^{K\times K}$:
$\Delta A=f_{\theta}(A_{prior},F_{s},F_{s}^{k})$ (16)
where $A_{prior}\in\{0,1\}^{K\times K}$ is the unweighted graph input provided by the user (like in GraphCape), and $F_{s}\in\mathbb{R}^{hw\times C}$ and $F_{s}^{k}\in\mathbb{R}^{K\times C}$ are the support image and keypoint features.
To handle the few-shot setting, our pose-graph predictor refines the features of each support image independently, allowing the model to account for structure on a per-image basis. After refinement, we average the resulting structure features across all support images and predict a unified adjacency matrix. This aggregation enables the method to exploit structural cues from all available support instances, ensuring that the final structure estimate benefits from as much visible information as possible. Note that the prior skeletal relations ($A_{prior}$) are category-specific and therefore identical across all support images.
C.3 Markov Attention Bias
In category-agnostic pose estimation, encoding spatial relations between keypoints is beneficial for robust localization, especially when dealing with novel objects (Hirschorn and Avidan, 2024; Rusanovsky et al., 2024). Transformer models naturally have a global receptive field, allowing each node (or token) to attend to all others within a layer. However, this flexibility introduces a challenge: Transformers lack inherent structural constraints, so positional dependencies must be explicitly encoded to reflect local relations. While this is often done in sequence data through absolute or relative positional encodings, graphs pose a different challenge, as nodes are not arranged linearly and connectivity is defined by edges.
Our approach aims to better utilize the structural dependencies between keypoints for CAPE. We build on the foundations set by GraphCape (Hirschorn and Avidan, 2024), which incorporates GCN layers into the feed-forward layers to propagate structural information. However, the GCN layers used in that model were limited by their fixed, nearest-neighbor receptive field, which restricts the model’s ability to capture more complex or distant connections between keypoints.
Thus, we further integrate the graph-prior into the architecture. We follow Ying et al.(Ying et al., 2021), adding a bias term based on graph connectivity to the self-attention mechanism in the decoder. Denote $a_{ij}$ as the $(i,j)$-element of the Query-Key product attention matrix $a$, resulting in:
@@ C.3 Markov Attention Bias
This bias term is highly effective in capturing general structure, boosting the performance of GraphCape by around 0.5%. However, it assumes a discrete distance between nodes and is applicable for unweighted adjacency matrices. Our primary objective is to predict real-valued edges, representing the strength of the structural keypoints connections. Thus, inspired by Ma et al.(Ma et al., 2023a), we treat the normalized adjacency matrix $\widetilde{A}$ as a right stochastic matrix.
A stochastic matrix is commonly used to describe the transitions in a Markov chain, where each element $(\widetilde{A})_{ij}$ represents the probability of moving from one state (or node) $v_{i}$ to another state $v_{j}$ in a single step. Specifically, the entry $(\widetilde{A}^{k})_{ij}$ gives the probability of transitioning from node $v_{i}$ to node $v_{j}$ in exactly $k$-hops. This process allows us to capture both direct and indirect relations between keypoints, enabling the model to consider more distant keypoints that may influence localization. Thus, we build the following matrix:
@@
where $I$ is the identity matrix and the parameter $K$ controls the maximum number of hops considered. We incorporate this graph characteristic into our model’s attention mechanism to enable more nuanced and flexible structural priors. The complete decoder layer is illustrated in Figure 16. Based on Equations 17 and 18, we use the following bias term:
@@
Figure 16: Graph Decoder Prediction Layer. Overview of the Transformer decoder architecture, adapted from the GraphCape design. The decoder consists of self-attention, cross-attention, and a graph-based feed-forward network. We incorporate a Markov Attention Bias into the self-attention mechanism to encourage structural keypoint interactions. Self-attention facilitates adaptive interactions among support keypoints, while cross-attention extracts localization information from the input features. Finally, the decoder refines keypoint features and outputs location predictions.
C.4 Training Scheme
@@
We begin by training the base model and establishing strong foundational features necessary for robust localization. Optimization is done using $\mathcal{L}_{offset}$ which is the $L_{1}$ localization loss as in (Shi et al., 2023b). Once the model has converged, we integrate the skeleton predictor. This stage allows the model to further adapt by incorporating specific structural insights provided by the skeleton predictor. For this phase, we add the adjacency loss, resulting in:
@@
In the final phase, we maintain the frozen feature extractor and freeze the skeleton predictor. Then, we integrate the Markov bias attention. This final stage allows the model to strengthen its capacity to interpret spatial dependencies between keypoints.
@@
C.5 Implementation Details
For a fair comparison, training parameters, data augmentations, and data pre-processing are kept the same as in previous works. In addition, the backbone size was matched to other works, thus, we use the smallest ($\sim$20M parameters) versions of SwinV2 and DinoV2.
The graph predicting network, encoder, and keypoint prediction decoder have 3 layers. For the Markov Bias Attention, we use a maximum of $K=4$ hops, and for $\mathcal{L}_{adj}$ we mask 50% of keypoints, using $\lambda_{adj}=1$. The model is built upon MMPose framework (Contributors, 2020), trained using Adam optimizer with a batch size of 16, the learning rate is $10^{-5}$, and decays by 10× on the 160th and 180th epoch. Each phase is trained for 100 epochs. Training takes $\sim$ 8 hours on a single Nvidia A100 GPU.
Evaluation of SCAPE (Liang et al., 2024) and FMMP (Chen et al., 2025a) was done by removing the keypoint identifiers (which were shown to be inapplicable for real-world scenarios (Nguyen et al., 2024; Yang et al., 2024; Hirschorn and Avidan, 2024)) and training a network using their official code. As FMMP leverages a larger number of keypoints, it experienced a relatively greater performance drop without these identifiers. For SCAPE, we use DinoV2 as the backbone to ensure a fair comparison. And, since FMMP relies on multi-scale features, we adopt a modern pre-trained multi-scale vision transformer backbone (PVT-V2). We also used GraphCape and CapeFormer official code for evaluation and visualizations.